Circular No. 6386

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995N IN MCG –2-38-017 S. D. Van Dyk, University of California at Berkeley; R. A. Sramek,

National Radio Astronomy Observatory (NRAO); K. W. Weiler and M. J. Montes, Naval Research Laboratory; and N. Panagia, ESA/Space Telescope Science Institute, report: “New observations of the peculiar type-II supernova 1995N, made on 1995 June 16.19 UT with the NRAO’s Very Large Array, show detection at wavelength 3.6 cm (8.4 GHz) of a source with a flux density of 3.90 ± 0.21 mJy at α = 14^{h}49^{m}28^{s}.313, δ = −10^{o}10′13′′.92 (equinox J2000,0; ± 0′′.2 in each coordinate). This is coincident to within 1′′ with an optical position for SN 1995N (position end figures 28^{s}.29, 13′′.7) measured by Van Dyk from an optical R-band image made on 1995 May 23 at the Lick Observatory 1.0-m telescope; note a slight discrepancy with the optical position reported on IAUC 6170 (position end figures 28^{s}.27, 15′′.4). The source was again detected at this wavelength on 1995 Sept. 15.96 with a flux density of 3.98 ± 0.22 mJy. If SN 1995N behaves as the peculiar type-II SNe 1988Z (Van Dyk et al. 1993, Ap.J. 419, L69) and 1986J (Weiler et al. 1990, Ap.J. 364, 611), we expect the flux density to increase slowly at centimeter wavelengths and eventually reach ∼ 20 mJy, ∼ 2000 times the centimeter radio luminosity of Cas A for an assumed distance of ∼ 24 Mpc. Radio monitoring is continuing at this and other wavelengths.”

COMET C/1995 O1 (HALE-BOPP) N. Biver, Observatoire de Paris-Meudon (OPM); D. Ballantyne, J. K.

Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo; and D. Bockelée-Morvan, J. Crovisier and H. Rauer, OPM, report the detection of methanol in comet C/1995 O1 with the James Clerk Maxwell Telescope: “From the sum of spectra observed on Mar. 16 and Apr. 5 and 6, the (4,1)–(4,0) A line at 307 GHz and the (2,1)–(2,0) A line at 304 GHz have integrated intensities in the antenna temperature scale of 35 ± 5 and 43 ± 6 mK km/s, respectively. In the thermal equilibrium assumption, this corresponds to a temperature of ∼ 20 K and to Q(CH_{3}OH) = 9 × 10^{26} molecules/s. The average of the two lines has a width of 1.4 ± 0.2 km/s and is slightly blueshifted (−0.04 ± 0.04 km/s) with respect to the comet. This is similar to the line profile of HCN reported on IAUC 6377, but significantly broader than the lines of CO observed in the comet.”

1996 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green